# GalaxyGenius: A Mock Galaxy Image Generator for Various Telescopes from Hydrodynamical Simulations

### v1

@@ 3.4 JWST NIRCam
The James Webb Space Telescope (JWST, Sabelhaus & Decker (2004)) is a cutting-edge observatory designed to operate primarily in the infrared wavelength range, covering 0.6 to 28 microns. With a 6.5-meter segmented primary mirror, JWST provides a collecting area over six times larger than the HST, enabling unprecedented sensitivity and resolution for observations in the infrared.
@@
The PSF for each filter can be retrieved from STPSF 888https://stpsf.readthedocs.io/en/latest/ (Perrin et al., 2015). The background emission for JWST has been extensively studied in Rigby et al. (2023) and can be obtained from JWST Background Tool (JBT) 999https://github.com/spacetelescope/jwst_backgrounds. Given a pointing and observational date, the background emission including ISM, Zodiacal light, stray light and thermal emission can be calculated from 0.5 to 30 micron. We similarly take 600 s exposure and derive the background and instrumental noise for each band of JWST. Here, we take F070W, F150W, F200W, and F182M from short and F356W and F444W from long-wavelength channels respectively and display bandpass images in Figure 6.
@@ 3.5 Roman
The Nancy Grace Roman Space Telescope (Roman, Spergel et al. (2015)), formerly known as WFIRST (Wide Field Infrared Survey Telescope), is a next-generation 2.36 m space-observatory developed by NASA. Scheduled for launch in the mid-2020s, Roman is designed to address key astrophysical questions, including the nature of dark energy and dark matter, the detection of exoplanets, the formation and evolution of galaxies.
The Wide Field Instrument (WFI) on Roman operates primarily in the infrared wavelength range, covering 0.48 to 2.3 microns, employ seven photometric filters and a grism for slitless spectroscopy. The pixel scales of each photometric filter is . The exposure time is similarly considered to be 600 s. We take the maximum value of dark current and readout noise demonstrated in the Roman user documentation 101010https://roman-docs.stsci.edu/roman-instruments-home for calculation of instrumental and background noises, while the PSFs are obtained from STPSF for Roman given the name of filter. We illustrate the mock images in F062, F087, F106, F129, F158, F184 and F213 bands in Roman respectively in Figure 7.
@@ 3.6 HSC
The Hyper Suprime-Cam Subaru Strategic Program (HSC-SSP, Aihara et al. (2018)) is a premier imaging survey conducted using the Hyper Suprime-Cam (HSC) on the 8.2-meter Subaru Telescope in Hawaii. It leverages HSC’s 1.5-degree field of view and exceptional resolution with a median seeing arcseconds to achieve deep, wide-field imaging across three tiers: Wide (1,400 square degrees, ), Deep (26 square degrees, ), and Ultra-Deep (3.5 square degres, ). The survey employs multi-band photometry in five broad bands ( and ) and speciallized narrow-band filters, enabling precise photometric redshift estimations and the detection of emission line galaxies. Built on the LSST software framework, the HSC-SSP pipeline handles large-scale data reduction, providing calibrated images, coadded mosaics, and comprehensive catalogs (Bosch et al., 2018). Scientifically, HSC-SSP addresses a wide range of objectives, including mapping dark matter via weak lening, studying galaxy evolution, and probing the early universe with high-redshift galaxy detections, making it a cornerstone for cosmological and astrophysical research.
@@ 3.7 Pseudo-color images
To visualize the sample outcomes of GalaxyGenius (see Figure 9), we generated pseudo-color images of selected spiral galaxies for CSST, Euclid, HST, JWST, Roman, and HSC, using the make_rgb routine in Astropy 131313https://docs.astropy.org/en/stable/visualization/rgb.html# based on the data products described in the previous subsections. However, it is important to note that optimal parameters for creating pseudo-color images require evaluation across a wide field for various observations, while we employ the default settings 141414https://docs.astropy.org/en/stable/visualization/rgb.html#rgb-images-using-arbitrary-stretching for the purpose of demonstration.
@@
To further illustrate the resolving power of the aforementioned instruments, we select another spiral galaxy with a subhaloID of 11 at (snapshot 84). The corresponding pseudo-color images can be created with similar processes to those for generating the images in the panel (a) in Figure 9, but the sizes of cutouts differ due to the angular size of the mock images, which are for the face-on case (upper row) and for the edge-on case (lower row), respectively. As displayed in the panel (b), the effects of PSF smearing are much more pronounced than the earlier galaxy at , highlighting the significance of using GalaxyGenius for the investigation of galaxy morphologies in more distant galaxies observed with space-borne telescopes.
@@ 3.8 EAGLE simulation
Galaxy images generated utill now are from IllustrisTNG simulations. However, GalaxyGenius is not constraint to specific hydrodynamical simulations. To demonstrate the capability of generating galaxies for general simulation, we employ EAGLE simulation as a case study. Two galaxies with GalaxyIDs of 21109761 and 20595120 at from RefL0100N1504 simulation in the EAGLE suite are considered. The parameters for these galaxies required by interface of data preprocessing module are acquired from SQL system in EAGLE’s official website 151515https://icc.dur.ac.uk/Eagle/database.php. Stars, star-forming regions, and dust elements and their corresponding properties outlined in Table 1 are retrieved from snapshot data using pyread_eagle 161616https://github.com/kyleaoman/pyread_eagle. It is important to note that, unlike TNG simulations, the smoothing length for gas particles should also be extracted. This is because EAGLE simulations are conducted using a Smoothed Particle Hydrodynamics (SPH) method known as Gadget-2 (Springel, 2005) for both stars and gases.
Using these inputs, the preprocessing module creates the execution file for SKIRT. Once this step is complete, the subsequent procedures are identical to those used for IllustrisTNG simulations. Galaxy images anticipated by CSST and JWST are generated and displayed in Figure 10. We notice that they are elliptical and spiral galaxies respectively, consistent with results from EAGLE. In contrast to the sample images presented in Section 3, Figure 3.8 exhibits several companion sources around the central image. This phenomenon arises from the fact that, unlike the TNG dataset, which stores its particle data based on individual subhalos, the particles from EAGLE are retrieved in a box based on central coordinate of a specific subhalo and boxlength mentioned in Section 2.2. Consequently, particles from companion sources are also included. However, the companion sources may have some issues due to incomplete extractions limited by boxlength. Anyway, above analysis highlight the GalaxyGenius’s adaptability to different simulations and its ability to produce consistent results across varied datasets.
@@ 4 Discussion
In this section, we firstly make several discussions on dust recipes and dust models. And then we explain some limitations of generating galaxies from hydrodynamical simulations using SKIRT project.
@@ 4.1 Dust
Firstly, we investigate the impact of two dust recipes mentioned in Section 2.2 on galaxy generation. Figure 11 presents a scatter plot for gas density versus temperature within selected box of Subhalo 31 and 253881. The colorbar indicates the gas mases. The densities are directly retrieved from TNG and are normalized by a threshold that distinguishes between star-formation and non-star-formation. The temperatures are calculated using Equation 3. The two recipes Camps et al. (2016) and Torrey et al. (2012) are indicated by green and red dashed lines, respectively.
@@
We observe that the recipe proposed by Torrey et al. (2012) tends to incorporate a higher number of dust particles. The precise numbers are presented in Table 2. Ideal near-infrared SEDs generated by SKIRT for both subhalos are displayed in Figure 12, where the lower panels display the normalized residuals. From this figure, we notice that the fluxes of Subhalo 31 exhibits a decrease in lower wavelength range when employing the recipe by Torrey et al. (2012). For Subhalo 253881, the SEDs at face-on direction are identical regardless of the recipe, while edge-on SEDs all experience a decrease at similar wavelength range. Additionally, the flux decrease is more prominent for Torrey et al. (2012) recipe, due to more dust introduced. The observed phenomenon of Subhalo 253881 at face-on direction can be potentially attributed to the substantial number of BC03 particles, leading to absence of any impact from additional dust on star-forming regions. Figure 13 depicts the spatial distribution of dust and star-forming regions for both dust recipes, revealing that the recipe given by Torrey et al. (2012) covers a larger area compared to the one given by Camps et al. (2016), which helps explain the SED behavior of Subhalo 253881 observed in Figure 12. Additionally, as is shown in the right panel, for spiral galaxies like Subhalo 253881, the Torrey et al. (2012) recipe provides a more complete dust distribution around the disk plane.
@@
On the other hand, we explore the influence of dust models that SKIRT project incorporated, including Zubko (Zubko et al., 2004), DraineLi (Draine & Li, 2007), and Themis (Jones et al., 2017). Themis comprises two families of dust particles: amorphous silicates and amorphous hydrocarbons, while Zubko and DraineLi also include polycyclic aromatic hydrocarbons (PAH) grains. The number of grain size bins employed to discretize the thermal emission calculations can be customized. In this study, we adopt 15 grain size bins for each dust mixture population, as recommended by Kapoor et al. (2021). The impact of employing distinct dust models on the SEDs is depicted in Figure 14. We notice that dust models indeed modify the flux curve for highly star-forming galaxies, which are predominantly composed of star-forming regions. While for galaxies with limited star-formation like Subhalo 253881, this effect is negligible in considered wavelength range.
@@ 4.2 Limitations
Current hydrodynamical simulations have significantly advanced our understanding of galaxy formation and evolution by modeling large cosmological volumes with high resolution. However, they are subject to several limitations. One major constraint is spatial and mass resolution (Pillepich et al., 2018), which affects the ability to resolve small-scale structure such as individual stars or molecular clouds. Limited resolution can lead to oversimplifed representations of physical processes, such as star formation and feedback mechanisms, which are often implemented via subgrid models rather than being explicitly resolved. Additionally, the assumptions in subgrid physics, such as recipes for stellar feedback, black hole accretion, and gas cooling, can introduce uncertainties and biases in simulation outcomes (Matsumoto et al., 2023). Another significant challenge is the limited volume of simulations, which restricts their ability to fully capture the large-scale structures of the universe. This limitations also results in a limited number of subhalos available for generating galaxies. For instance, in snapshot-94 of TNG100, which we use as an example in Section 2.2, there are approximately 6,000 galaxies with stellar masses exceeding . While one potential workaround is to increase the number of galaxies by considering multiple viewing angles, this approach does not address the lack of diversity among the galaxies. Galaxies generated from the same subhalo share identical intrinsic properties, such as stellar mass, star formation rate (SFR), and other characteristics, limiting the ability to explore a broader range of galactic properties and behaviors.
@@
On the other hand, SKIRT radiative transfer project is a powerful tool to simulate the interaction of radiation with dust and gas in astrophysical environments. Nonetheless, it has its own set of limitations. SKIRT simulations, especially those involving complex geometries or high-resolution grids, can be computationally demanding, requiring substantial memory and processing power, which can restrict the feasibility of highly detailed simulations. In regions with extremely high optical depths, accurately modeling radiative transfer becomes challenging (Krieger & Wolf, 2023; Camps & Baes, 2018; Baes et al., 2019), as photon packets may undergo numerous interactions, leading to convergence issues and increased computational demands. As with all Monte Carlo methods, SKIRT simulations are subject to statistical noise, particularly in regions with low photon packet counts; mitigating this noise requires launching a large number of photon packets, which further increases computational load. Moreover, the SED families within SKIRT, such as BC03, FSPS and MAPPINGS-III are usually employed at low and intermediate redshifts. Their applicability to high redshift galaxies remains an open issue. Studies indicate that a single SSP model may not fully capture the complexities of early-type galaxies at high redshifts (López-Corredoira et al., 2017), potentially affecting the realism of the generated SEDs and galaxy images. Additionally, MAPPINGS-III requires three free parameters as shown in Table 1, and their values need to be determined through calibrations with real observations.
@@
These limitations underscore the importance of ongoing advancements in finer and larger hydrodynamical simulations, and the development of stellar populations at high redshift. By addressing these challenges, the astrophysical community can further enhance the accuracy and predictive power of these tools, enabling deeper insights into the complex processes shaping galaxies across the universe, thus improving the realism and fidelity for simulated galaxies from hydrodynamical simulations.
@@
Furthermore, in this work, we primarily focus on generating galaxy images. In addition to bandpass images, with data cubes and SEDs, simulating observations by IFU (Harborne, 2019; Wang et al., 2024; Sarmiento et al., 2023; Bottrell & Hani, 2022; Nanni et al., 2023; Çakır et al., 2024), slit spectrograph (Fagioli et al., 2018; Kirkby et al., 2021) and slitless spectrograph (Kümmel et al., 2009; Neveu et al., 2024; Taylor, 2020; Brammer et al., 2012; Wen et al., 2024; Zhou et al., 2024) are also possible and straightforward by implementing external functions for post-processing.

### v2

@@ 3.4 JWST NIRCam
The James Webb Space Telescope (JWST, Sabelhaus & Decker (2004)) is a cutting-edge observatory designed to operate primarily in the IR wavelength range, covering 0.6 to 28 microns. With a 6.5-meter segmented primary mirror, JWST provides a collecting area over six times larger than the HST, enabling unprecedented sensitivity and resolution for observations in the IR.
@@
The PSF for each filter can be retrieved from STPSF 888https://stpsf.readthedocs.io/en/latest/ (Perrin et al., 2015). The background emission for JWST has been extensively studied in Rigby et al. (2023) and can be obtained from JWST Background Tool (JBT) 999https://github.com/spacetelescope/jwst_backgrounds. Given a pointing and observational date, the background emission including ISM, Zodiacal light, stray light, and thermal emission can be calculated from 0.5 to 30 micron. We similarly took a 600 s exposure and derive the background and instrumental noise for each band of JWST. We took F070W, F150W, F200W, and F182M from the short-wavelength and F356W and F444W from long-wavelength channels, respectively. We display bandpass images in Figure 6.
@@ 3.5 Roman
The Nancy Grace Roman Space Telescope (Roman, Spergel et al. (2015)), formerly known as WFIRST (Wide Field Infrared Survey Telescope), is a next-generation 2.36 m space-observatory developed by NASA. Scheduled for launch in the mid-2020s, Roman is designed to address key astrophysical questions, including the nature of dark energy and dark matter, the detection of exoplanets, and the formation and evolution of galaxies.
The Wide Field Instrument (WFI) on Roman operates primarily in the IR wavelength range, covering 0.48 to 2.3 microns, employ seven photometric filters and a grism for slitless spectroscopy. The pixel scales of each photometric filter is . The exposure time is similarly considered to be 600 s. We took the maximum value of dark current and readout noise demonstrated in the Roman user documentation 101010https://roman-docs.stsci.edu/roman-instruments-home for the calculation of instrumental and background noises, while the PSFs were obtained from STPSF for Roman given the name of filter. We illustrate the mock images in F062, F087, F106, F129, F158, F184, and F213 bands in Roman, respectively, in Figure 7.
@@ 3.6 HSC
The Hyper Suprime-Cam Subaru Strategic Program (HSC-SSP, Aihara et al. (2018)) is a premier imaging survey conducted using the Hyper Suprime-Cam (HSC) on the 8.2-meter Subaru Telescope in Hawaii. It leverages HSC’s 1.5-degree field of view and exceptional resolution with a median seeing arcseconds to achieve deep, wide-field imaging across three tiers: wide (1,400 square degrees, ), deep (26 square degrees, ), and ultra-deep (3.5 square degres, ). The survey employs multi-band photometry in five broad bands (, and ) and specialized narrowband filters, enabling precise photometric redshift estimations and the detection of emission line galaxies. Built on the LSST software framework, the HSC-SSP pipeline handles large-scale data reduction, providing calibrated images, coadded mosaics, and comprehensive catalogs (Bosch et al., 2018). Scientifically, HSC-SSP addresses a wide range of objectives, including mapping dark matter via weak lensing, studying galaxy evolution, and probing the early universe with high-redshift galaxy detections, making it a cornerstone for cosmological and astrophysical research.
@@ 3.7 Pseudo-color images
To visualize the sample outcomes of GalaxyGenius, we generated pseudo-color images of selected spiral galaxies for CSST, Euclid, HST, JWST, Roman, and HSC, using the make_rgb routine in Astropy 131313https://docs.astropy.org/en/stable/visualization/rgb.html# based on the data products described in the previous subsections. However, it is important to note that optimal parameters for creating pseudo-color images require an evaluation across a wide field for various observations, while we employed the default settings 141414https://docs.astropy.org/en/stable/visualization/rgb.html#rgb-images-using-arbitrary-stretching for the purpose of demonstration.
@@
To further illustrate the resolving power of the aforementioned instruments, we selected another spiral galaxy with a subhaloID of 11 at (snapshot 84). The corresponding pseudo-color images can be created with similar processes to those for generating the images in the panel (a) in Figure 9, but the sizes of cutouts differ due to the angular size of the mock images, which are for the face-on case (upper row) and for the edge-on case (lower row), respectively. As displayed in the panel (b), the effects of PSF smearing are much more pronounced than the galaxy at , highlighting the significance of using GalaxyGenius for the investigation of galaxy morphologies in more distant galaxies observed with space-borne telescopes.
@@ 3.8 EAGLE simulation
Galaxy images generated utill now are from IllustrisTNG simulations. However, GalaxyGenius is not constraint to specific hydrodynamical simulations. To demonstrate the capability of generating galaxies for general simulation, we employed EAGLE simulation as a case study. Two galaxies with GalaxyIDs of 21109761 and 20595120 at from RefL0100N1504 simulation in the EAGLE suite are considered. The parameters for these galaxies required by interface of data preprocessing module are acquired from SQL system in EAGLE’s official website 151515https://icc.dur.ac.uk/Eagle/database.php. Stars, star-forming regions, and dust elements and their corresponding properties outlined in Table 1 were retrieved from snapshot data using pyread_eagle 161616https://github.com/kyleaoman/pyread_eagle. It is important to note that unlike TNG simulations, the smoothing length for gas particles should also be extracted. This is because EAGLE simulations are typically conducted using a smoothed particle hydrodynamics (SPH) method known as Gadget-2 (Springel, 2005) for both stars and gases.
Using these inputs, the preprocessing module creates the execution file for SKIRT. Once this step is complete, the subsequent procedures are identical to those used for IllustrisTNG simulations. Galaxy images anticipated by CSST and JWST are generated and displayed in Figure 10. We notice that they are elliptical and spiral galaxies, respectively, consistent with results from EAGLE. In contrast to the sample images presented in Section 3, Figure 3.8 exhibits several companion sources around the central image. This phenomenon arises from the fact that unlike the TNG dataset, which stores its particle data based on individual subhalos, the particles from EAGLE are retrieved in a box based on central coordinate of a specific subhalo and box length (as mentioned in Section 2.2). Consequently, particles from companion sources are also included. However, the companion sources may have some issues due to incomplete extractions limited by box length. Anyway, the above analyses highlight the GalaxyGenius’s adaptability to different simulations and its ability to produce consistent results across varied datasets.
@@ 4 Discussion
In this section, we first offer several discussions on dust recipes and dust models. Then we explain some limitations of generating galaxies from hydrodynamical simulations using the SKIRT project.
@@ 4.1 Dust
First, we investigated the impact of two dust recipes mentioned in Section 2.2 on galaxy generation. Figure 11 presents a scatter plot for gas density versus temperature within selected box of subhalo 31 and 253881. The colorbar indicates the gas mases. The densities were directly retrieved from TNG and normalized by a threshold that distinguishes between star formation and non-star formation. The temperatures are calculated using Equation 3. The two recipes Camps et al. (2016) and Torrey et al. (2012) are indicated by green and red dashed lines, respectively.
@@
We observe that the recipe proposed by Torrey et al. (2012) tends to incorporate a higher number of dust particles. The precise numbers are presented in Table 2. Ideal NIR SEDs generated by SKIRT for both subhalos are displayed in Figure 12, where the lower panels display the normalized residuals. From this figure, we notice that the fluxes of subhalo 31 exhibits a decrease in lower wavelength range when employing the recipe by Torrey et al. (2012). For subhalo 253881, the SEDs at face-on direction are identical regardless of the recipe, while edge-on SEDs all experience a decrease at similar wavelength range. Additionally, the flux decrease is more prominent for Torrey et al. (2012) recipe, due to more dust introduced. The observed phenomenon of subhalo 253881 at face-on direction can be potentially attributed to the substantial number of BC03 particles, leading to absence of any impact from additional dust on star-forming regions. Figure 13 depicts the spatial distribution of dust and star-forming regions for both dust recipes, revealing that the recipe given by Torrey et al. (2012) covers a larger area compared to the one given by Camps et al. (2016), which helps explain the SED behavior of subhalo 253881 observed in Figure 12. Additionally, as is shown in the right panel, for spiral galaxies such as subhalo 253881, the Torrey et al. (2012) recipe provides a more complete dust distribution around the disk plane.
@@
On the other hand, we explore the influence of dust models that SKIRT project incorporated, including Zubko (Zubko et al., 2004), DraineLi (Draine & Li, 2007), and Themis (Jones et al., 2017). Themis comprises two families of dust particles: amorphous silicates and amorphous hydrocarbons, while Zubko and DraineLi also include polycyclic aromatic hydrocarbons (PAH) grains. The number of grain size bins employed to discretize the thermal emission calculations can be customized. In this study, we adopt 15 grain size bins for each dust mixture population, as recommended by Kapoor et al. (2021). The impact of employing distinct dust models on the SEDs is depicted in Figure 14. We notice that dust models indeed modify the flux curve for highly star-forming galaxies, which are predominantly composed of star-forming regions. While for galaxies with limited star-formation such as subhalo 253881, this effect is negligible in considered wavelength range.
Additionally, this study only implement the dust attenuation effect from the dust particles inside galaxies themselves. A rigorous mock observation may also necessitate the dust extinction effect along the line-of-sight, i.e. the dust within the IGM and the Milky Way (Thomas et al., 2017, 2020, 2021; Cardelli et al., 1989; Fitzpatrick, 1999; Sofue & Nakanishi, 2016; Misiriotis et al., 2006). However, it is generally believed that the dust extinction effect from the IGM is not significant, where the observational constraint is also very limited (Xie et al., 2015). The foreground Milky Way dust extinction mainly depends on the galactic latitude of the target, which has been extensively studied and is easily to calibrate (Schlegel et al., 1998; Li et al., 2018; Zhang et al., 2023; Zhang & Green, 2025). Users can implement their own modules to include and calibrate these effects.
@@ 4.2 Limitations
Current hydrodynamical simulations have significantly advanced our understanding of galaxy formation and evolution by modeling large cosmological volumes with high resolution. However, they are subject to several limitations. One major constraint is spatial and mass resolution (Pillepich et al., 2018), which affects the ability to resolve small-scale structure such as individual stars or molecular clouds. Limited resolution can lead to oversimplifed representations of physical processes, such as star formation and feedback mechanisms, which are often implemented via subgrid models, rather than being explicitly resolved. Additionally, the assumptions in subgrid physics, such as recipes for stellar feedback, black hole accretion, and gas cooling, can introduce uncertainties and biases in simulation outcomes (Matsumoto et al., 2023). Another significant challenge is the limited volume of simulations, which restricts their ability to fully capture the large-scale structures of the universe. This limitations also results in a limited number of subhalos available for generating galaxies. For instance, in snapshot-94 of TNG100 (which we use as an example in Section 2.2), there are approximately 6,000 galaxies with stellar masses exceeding . While one potential workaround is to increase the number of galaxies by considering multiple viewing angles, this approach does not address the lack of diversity among the galaxies. Galaxies generated from the same subhalo share identical intrinsic properties, such as stellar mass, SFR, and other characteristics, limiting the ability to explore a broader range of galactic properties and behaviors.
@@
On the other hand, SKIRT radiative transfer project is a powerful tool to simulate the interaction of radiation with dust and gas in astrophysical environments. Nonetheless, it has its own set of limitations. SKIRT simulations, especially those involving complex geometries or high-resolution grids, can be computationally demanding, requiring substantial memory and processing power, which can restrict the feasibility of highly detailed simulations. In regions with extremely high optical depths, accurately modeling radiative transfer becomes challenging (Krieger & Wolf, 2023; Camps & Baes, 2018; Baes et al., 2019), as photon packets may undergo numerous interactions, leading to convergence issues and increased computational demands. As with all Monte Carlo methods, SKIRT simulations are subject to statistical noise, particularly in regions with low photon packet counts; mitigating this noise requires launching a large number of photon packets, which further increases computational load. Moreover, the SED families within SKIRT, such as BC03, FSPS, and MAPPINGS-III are usually employed at low and intermediate redshifts. Their applicability to high redshift galaxies remains an open issue. Studies indicate that a single SSP model may not fully capture the complexities of early-type galaxies at high redshifts (López-Corredoira et al., 2017), potentially affecting the realism of the generated SEDs and galaxy images. Additionally, MAPPINGS-III requires three free parameters as shown in Table 1 and their values need to be determined through calibrations with real observations.
@@
These limitations underscore the importance of ongoing advancements in finer and larger hydrodynamical simulations, and the development of stellar populations at high redshift. By addressing these challenges, the astrophysical community can further enhance the accuracy and predictive power of these tools, enabling deeper insights into the complex processes shaping galaxies across the universe, thereby improving the realism and fidelity for simulated galaxies from hydrodynamical simulations.
@@
Furthermore, in this work, we are primarily focused on generating galaxy images. In addition to bandpass images, with data cubes and SEDs, it is also possible to simulate observations by IFU (Harborne, 2019; Wang et al., 2024; Sarmiento et al., 2023; Bottrell & Hani, 2022; Nanni et al., 2023; Çakır et al., 2024), slit spectrograph (Fagioli et al., 2018; Kirkby et al., 2021), and slitless spectrograph (Kümmel et al., 2009; Neveu et al., 2024; Taylor, 2020; Brammer et al., 2012; Wen et al., 2024; Zhou et al., 2024) in a straightforward way by implementing external functions for post-processing.